# Application of the FRADO model of BLR formation to the Seyfert galaxy NGC 5548 and the first step toward determining the Hubble constant

### v2

@@ 3.3 BLR structure and reprocessing
To introduce the contribution from the BLR, we first describe its properties. In this study, we do not parameterize the structure of the BLR using arbitrary numbers or functions; instead, we calculate it based on the FRADO model, which was qualitatively proposed by Czerny & Hryniewicz (2011) and has been further developed in several subsequent studies (Czerny et al., 2017; Naddaf et al., 2021; Naddaf & Czerny, 2022; Naddaf et al., 2023; Naddaf & Czerny, 2024). Specifically, we utilize the code from Naddaf et al. (2021), which provides a detailed numerical description of the wavelength-dependent cross-section for dusty particles. According to this model, radiatively dust-driven pressure lifts the clouds from the disk surface, while preserving their angular momentum derived from the Keplerian motion of the disk surface. As a cloud is lifted, it becomes increasingly illuminated by the inner parts of the disk. However, if the cloud becomes too hot, the dust evaporates, allowing the cloud to continue its motion along a ballistic orbit. The global parameters of the model, such as black hole mass, Eddington rate, and metallicity, govern the behavior of the BLR. For lower black hole masses, lower Eddington ratios, and lower metallicities, the clouds form a failed wind, while in the opposite case, a fraction of the clouds may form an escaping wind. Hence, we emphasize that the inner and outer radii of the BLR, along with the statistical distribution of the clouds, are governed by global parameters such as the black hole mass and the Eddington ratio.
@@
where is the distance of a cloud from the center. This scaling of the emissivity, which inversely depends on the square of the distance, neglects the local efficiency of converting incident flux into BLR spectrum. However, since our model does not yet predict the local density, this simplification is reasonable at this stage of development. Later in the model, we fix the local density of the cloud at cm-3. For convenience, the value 4900 serves as a reference.
@@ 3.4 Photoionization calculations using CLOUDY
To determine the BLR spectral shape, we perform photoionization calculations using the code CLOUDY, version C23.00 (Chatzikos et al., 2023). In these computations, we simplify the 3-D cloud distribution by replacing the entire ensemble of clouds with a single representative cloud, positioned at the radius inferred from the time delay. For the analysis at this established distance, we adopt an incident radiation luminosity of (erg/s) = 44, a BLR distance from the central source of (cm) = 16, a constant hydrogen gas density of (cm-3) = 11, and a column density of (cm-2) = 23.5, following the prescriptions of (Korista & Goad, 2001; Panda et al., 2022). We assume solar metallicity and neglect turbulence in the medium.
@@
The shape of the incident SED is taken from the file ”NGC5548.sed” available in the CLOUDY database, which corresponds to the SED derived by Mehdipour et al. (2015). Their study estimated the starlight contribution to NGC 5548 using HST observations, providing a well-constrained incident SED that we adopt for our calculations. The resulting BLR emissivity profile, , is then incorporated into the final delay model. When we later account for the assumption of an unknown luminosity distance, the irradiating flux is adjusted accordingly to ensure consistency with the observed hard X-ray flux.
@@
Since CLOUDY does not compute higher-order Balmer lines, we incorporate a scaled component into the spectrum, following the method of Kovačević et al. (2014). This modification improves the spectral fit, as demonstrated in Pandey et al. (2024), though it only affects the wavelength range between 3646Å and 4000Å in the rest frame.
@@
Black Hole Mass | |||
Corona Height | 20 | 43 | 32 |
Inclination Angle | 40 | 45 | 40 |
Warm corona Inner Radius | - | 43 | 6 |
Warm Corona Temperature | - | ||
Warm corona optical depth/Photon index | - | 2.28 | 21.5 |
Inner Cold Disk Radius | 35 | 151 | 90 |
Outer Cold Disk Radius | 10000 | 282 | 10000 |
Color Correction | 2.4 | 1.0 | 1.0 |
BLR Contribution() | 12 | 40 | 36 |
Lamp Luminosity | |||
Eddington Ratio | 0.02 | 0.027 | 0.015 |
Note. Columns are (1) name of the parameters used in the model. Black hole mass in unit of , corona height in unit of , inclination angle in degree, inner disk radius and outer disk radius in , lamp luminosity in , (2) parameter values for Model A (3) parameter values for Model B, (4) parameter values for Model C. Warm corona optical depth or a photon index are alternative (equivalent) parameters of Comptonization models.
@@ 3.6 Combined time delay and NGC 5548 data fitting
To fit the data of NGC 5548, we utilize all the parameters listed in Table1, integrating the contributions from both the disk and the BLR as described in equation 7.
When searching for the optimal solution, we impose constraints from both the observed SED and the measured time delays. Specifically, we incorporate all 17 available time delay measurements and select 16 representative points, evenly distributed across the entire observed SED range. To systematically evaluate and compare different solutions, we integrate these constraints into a unified criterion:
where observed data, is the error in observed data, and value from the model.
@@ 4 Results
In this section we concentrate on testing the FRADO model against the spectral and time-delay data for NGC 5548. We fix the mass of the central black hole and the bolometric luminosity in our analysis. We assume that the distance to the source is known, and corresponds to the source redshift. The issue of the Hubble constant determination will be addressed separately in Section 5.
@@ 4.1 BLR properties from FRADO
The global model parameters adopted for NGC 5548, combined with the assumption of a dust sublimation temperature of 1500 K, allow us to uniquely determine the structure of the BLR. Notably, the FRADO model is specifically applicable to low-ionization lines such as H, Mg II, and Fe II. In this framework, the cloud positions within the BLR are determined by the FRADO model, with parameters relevant to NGC 5548: a black hole mass of , an Eddington ratio of 0.02, and an assumed metallicity 5 times solar.
@@
In this model, clouds are launched from the disk surface but remain relatively close to it, with the ratio of vertical height () to radial distance () staying below 5. This limited height is a direct consequence of the low Eddington ratio, giving the BLR a geometry that closely resembles a puffed-up disk surface. The ratio peaks near the inner edge of the BLR, which is defined by the dust sublimation temperature at approximately 2260 . At this radius, the Keplerian velocity is km s-1, implying that the full width at half maximum (FWHM) of the emission lines could reach up to twice this value. However, contributions from larger radii act to moderate this broadening.
@@
Next, we derive the BLR response function based on the cloud positions for an inclination angle of 40 degrees, which is likely representative of NGC 5548. The resulting shape, shown in Figure 2, exhibits two distinct peaks: one at shorter distances, around 2.6 days, corresponding to the clouds on the same side as the observer, and the other peak arising from the clouds on the opposite side of the black hole. The shape of our response function does not perfectly match the observed one in the data from Horne et al. (2021). Their projected response from the MEMECHO software for H (their Fig. 5, right panel) shows two peaks: one at days and the second one, much smaller, at days. Our first peak coincides quite well with their first peak but our second peak is located closer in, at days and it is relatively stronger than the second peak of Horne et al. (2021). However, it is noteworthy that no arbitrary parameters, such as inner or outer radii, were introducedonly global parameters were used. This lack of free parameters adds to the intrigue of the observed similarity between the modeled and real response functions.
@@
The measured time delay of the H line relative to the 5100 Å continuum during the 2014 campaign (January to June) was reported by Pei et al. (2017) as days, which is shorter than predicted by the radius–luminosity relation. This H delay ranks among the shortest ever recorded for NGC 5548. In contrast, the mean response function from the 19982001 campaign peaked around 20 days (Cackett & Horne, 2006). Horne et al. (2021) later explored the nature of this shorter delay in detail, finding that the response function exhibited a secondary peak at approximately days but began with a high value at zero time delay. Their observed map also supports the primary response originating from the near side of the BLR, facing the observer.
@@
Additionally, the model predicts an upper limit for the covering factor of the BLR. If the cloud distribution is not transparent, the region intercepts all radiation emitted within inclinations greater than the aspect ratio of the cloud distribution. As shown in Figure 1, the model suggests that less than of the radiation can be intercepted by the BLR, as the cloud height is generally small compared to the radius. This is notably lower than the to covering factor typically expected from BLRs (e.g. Baldwin et al., 1995; Korista & Goad, 2019; Panda, 2021; Panda et al., 2022). In contrast, the yearly variations observed in the line profile of NGC 5548 suggest that the disk can be distorted and/or precessing (e.g., Shapovalova et al., 2009, see their Fig. 2).
The distribution of BLR clouds enables us to roughly estimate the shape of the H line predicted by the model. In this estimation, we neglect the vertical cloud velocities and only consider the projected rotational velocities towards the observer. The resulting line shape, shown in Figure 3, reveals a two-peak structure that is even more pronounced than that observed in the data. The FWHM from the model approaches 8000 km s-1. In contrast, the measured FWHM for the H line in 2015 from the mean spectrum was km s-1 (Lu et al., 2022), while the FWHM measured from the RMS spectrum is slightly narrower, km s-1. Moreover, the FWHM values have varied in subsequent years, indicating the dynamic nature of the BLR.
@@ 4.2 The BLR emissivity
To model the emission characteristics of photoionized clouds, we perform CLOUDY simulations that account for both line and continuum contributions. During our computations, we consider both line and continuum emissions, incorporating radiation from both inward and outward directions to ensure equal visibility of the illuminated and un-illuminated sides of the clouds. Among the most prominent features that emerge are the Balmer and Paschen edges, as illustrated in Figure 4. The depths of these edges are governed by the specific model parameters adopted (see, e.g., the most recent study by Pandey et al., 2023).

### v3

@@ 3.3 BLR structure and reprocessing
To introduce the contribution from the BLR, we first describe its properties. In this study, we do not parameterize the structure of the BLR using arbitrary numbers or functions; instead, we calculate it based on the FRADO model, which was qualitatively proposed by Czerny & Hryniewicz (2011) and has been further developed in several subsequent studies (Czerny et al., 2017; Naddaf et al., 2021; Naddaf & Czerny, 2022; Naddaf et al., 2023; Naddaf & Czerny, 2024).
The basic idea behind the model is to associate the low-ionization parts of the BLR – responsible for emission lines, such as H, Mg II, and Fe II, with dust-driven massive winds. In stellar environments, such winds are typically denser, possess higher optical depths, and exhibit significantly slower outflow velocities. Similarly, in AGNs, dust-driven winds are expected to be launched from the outer disk, where the effective temperature falls below the dust sublimation threshold. This temperature constraint naturally defines the inner radius of the BLR. The nature of the wind depends on and Eddington ratio: for sufficiently massive black holes and high accretion rates, the wind escapes, contributing to the BLR outflow structure; otherwise, it forms a failed wind, in which the material ultimately falls back onto the disk. Additionally, irradiation from the central parts of the disk plays an increasingly important role in shaping the dynamics and ionization state of the wind as the cloud elevates above the disk plane.
Specifically, we utilize the code from Naddaf et al. (2021), which provides a detailed numerical description of the wavelength-dependent cross-section for dusty particles. According to this model, radiatively dust-driven pressure lifts the clouds from the disk surface, while preserving their angular momentum derived from the Keplerian motion of the disk surface. As a cloud is lifted, it becomes increasingly illuminated by the inner parts of the disk. However, if the cloud becomes too hot, the dust evaporates, allowing the cloud to continue its motion along a ballistic orbit. The global parameters of the model, such as , Eddington rate, and metallicity, govern the behavior of the BLR. For lower black hole masses, lower Eddington ratios, and lower metallicities, the clouds form a failed wind, while in the opposite case, a fraction of the clouds may form an escaping wind. Hence, we emphasize that the inner and outer radii of the BLR, along with the statistical distribution of the clouds, are governed by global parameters, such as and the Eddington ratio.
We assume a universal value of the dust sublimation temperature of 1500 K as representative for all grain species and sizes. It was suggested by Barvainis (1987) at the basis of the observational data for PG quasars. It is frequently used for zero order approximation, although it is well known that actually the sublimation temperature depends on the chemical composition as well as size of the dust grains (Draine & Lee, 1984; Baskin & Laor, 2018). However, calculating the range of evaporation temperature of each grain is time consuming and not appropriate for a pilot study. In addition, the chemical composition of grains in AGN is not firmly established, and the presence of the graphite with usual grain size distribution is highly unlikely (e.g. Czerny et al., 2004; Gaskell et al., 2004). We further discuss this issue in Section 6.3.
@@
where is the distance of a cloud from the center. This scaling of the emissivity, which inversely depends on the square of the distance, neglects the local efficiency of converting incident flux into BLR spectrum. However, since our model does not yet predict the local density, this simplification is reasonable at this stage of development. We fix the local density of the cloud at cm-3, and the column density at cm-2.
@@ 3.4 Photoionization calculations using CLOUDY
To determine the BLR spectral shape, we perform photoionization calculations using the code CLOUDY, version C23.00 (Chatzikos et al., 2023a). In these computations, we simplify the 3-D cloud distribution by replacing the entire ensemble of clouds with a single representative cloud, positioned at the radius inferred from the time delay. For the analysis at this established distance, we adopt an incident radiation luminosity of (erg/s) = 44, and a BLR distance from the central source of (cm) = 16 (Dalla Bontà et al., 2020), a constant hydrogen gas density of (cm-3) = 11, and a column density of (cm-2) = 23.5, following the prescriptions of (Korista & Goad, 2001; Panda et al., 2022). The adopted distance in our representative case corresponds to the measured time delay of the H line in NGC 5548. The assumed bolometric luminosity is also representative of this source, based on the redshift-inferred distance within the framework of standard cosmology. In this pilot study, we do not explore a wide range of gas densities or column densities in detail, as photoionization models, particularly single-zone models, still do not perfectly represent the full complexity of BLR properties (e.g. Netzer, 2020; Pandey et al., 2025; Floris et al., 2025). We assume a metallicity five times higher than solar, as this was favored by comparisons between the FRADO model and quasar spectra (Naddaf et al., 2021). Independent studies of quasar metallicity also support super-solar values (e.g. Śniegowska et al., 2021), and in the specific case of NGC 5548, high metallicity was advocated by Netzer (2020) based on line emissivity analysis. For simplicity, we neglect turbulence in the medium.
@@
The shape of the incident SED is taken from the file ”NGC5548.sed” available in the CLOUDY database, which corresponds to the SED derived by Mehdipour et al. (2015). Their study estimated the starlight contribution to NGC 5548 using HST observations, providing a well-constrained incident SED that we adopt for our calculations.
This is clearly an oversimplification, in two aspects. Each cloud receives the radiation of the same shape, only scaled down with the distance as specified in equation 7. This radiation includes both the hard X-ray radiation as well as the warm corona and the disk emission. Apart from the spectral shape, we also do not differentiate here between the arrival time from the hot corona, and warm extended corona. Accounting for such differences would require modeling the BLR response separately to the hard X-ray emission from the hot corona and to the emission from the warm corona, followed by merging several resulting transfer functions. As this is a pilot study, we have not undertaken that level of complexity at this stage, although it is feasible in principle, given that our spectral decomposition separates the contributions from all three components. For the current analysis, we adopt a fixed SED shape, a simplification that is unlikely to significantly affect our results. The hard X-ray power-law from the hot corona is consistent with SED, and the warm corona is also fitted in such way as to reproduce the spectrum. Slight error on the time delay introduced by the fact that some photons should come directly from the hot corona instead from the disk should not be essential, taking into account much larger size of the BLR than the disk.
The resulting BLR emissivity profile, , is then incorporated into the final delay model. When we later account for the assumption of an unknown luminosity distance, the irradiating flux is adjusted accordingly to ensure consistency with the observed hard X-ray flux.
@@
Since CLOUDY does not compute higher-order Balmer lines, we incorporate a scaled component into the spectrum, following the method of Kovačević et al. (2014). This modification improves the spectral fit, as demonstrated in Pandey et al. (2025), though it only affects the wavelength range between 3646Å and 4000Å in the rest-frame. The shape of this component is taken directly from Pandey et al. (2025); however, the emissivity still appears too low to fully capture the gradual decline observed beyond the Balmer edge in the data. More advanced modeling may be required to improve the fit in that region.
When we perform computations of the Model C for fixed luminosity distance the value of the incident flux in CLOUDY computations is fixed and corresponds to the observed luminosity and the time delay to the BLR as above. However, later on, in the Section 5 when the luminosity distance is treated as unknown, we scale the incident bolometric luminosity to the fitted accretion rate which is a function of the luminosity distance.
@@
| Black Hole Mass() | |||
| Corona Height( h) | |||
| Inclination Angle( i) | |||
| Warm corona Inner Radius | |||
| Warm Corona Temperature | |||
| Warm corona optical depth/Photon index | |||
| Inner Cold Disk Radius( ) | |||
| Outer Cold Disk Radius( ) | |||
| Color Correction | |||
| BLR Contribution() | 12 | 40 | 30 |
| Lamp Luminosity() | |||
| Eddington Ratio | |||
| Starlight |
Note. Columns are (1) name of the parameter used in the model, with values given in columns (2), (3) and (4) for Models A, B, and C, respectively. Black hole mass in unit of , corona height in unit of , inclination angle in degree, inner disk radius and outer disk radius in , lamp luminosity in . Warm corona optical depth or a photon index are alternative (equivalent) parameters of Comptonization models. Starlight normalization is given in erg s-1 cm-2 Å-1 at 5100 Å. References for the fixed parameters during fitting: (ref1) Netzer (2022); (ref2) Kammoun et al. (2021b); (ref3) Crenshaw et al. (2009); (ref4) Kubota & Done (2018); (a) fitted parameters. For model C we assume standard CDM cosmology, with km s-1 Mpc-1, .
@@ 3.6 Combined time delay and NGC 5548 data fitting
As mentioned in Section 3.4, we do not treat separately photons reaching the BLR directly from the hot corona and/or from the warm corona, and from the disk. Photons from the hot and warm corona are included in BLR computations as the incident continuum (i.e. they are included in the ), but their arrival time is the same as for the photon disks under consideration. This is done for the efficiency of computations when fitting the data. Replacing the incident continuum adopted for photoionization modeling with three separate spectral components would additionally require repeating photoionization computations at every computational step.
@@
To fit the data of NGC 5548, we utilize all the parameters listed in Table1, integrating the contributions from both the disk and the BLR as described in equation 3.6.
When searching for the optimal solution, we impose constraints from both the observed SED and the measured time delays. Specifically, we incorporate all 17 available time delay measurements and select 17 representative points, evenly distributed across the entire observed SED range. To systematically evaluate and compare different solutions, we integrate these constraints into a unified criterion:
where observed data, is the error in observed data, and value from the model. We assume the error in the spectral data of 7%, and the errors of the delay measurements are taken from Fausnaugh et al. (2016).
@@ 4 Results
In this section we concentrate on testing the FRADO model against the spectral and time-delay data for NGC 5548 and summarize the results obtained from our model fitting. We fix and the bolometric luminosity in our analysis. We assume that the distance to the source is known, and corresponds to the source redshift. The issue of the Hubble constant determination will be addressed separately in Section 5.
@@ 4.1 BLR properties from FRADO
The global model parameters adopted for NGC 5548, combined with the assumption of a dust sublimation temperature of 1500 K, allow us to uniquely determine the structure of the BLR. Notably, the FRADO model is specifically applicable to low-ionization lines such as H, Mg II, and Fe II. In this framework, the cloud positions within the BLR are determined by the FRADO model, with parameters relevant to NGC 5548: (Netzer, 2022), an Eddington ratio of 0.02 (Crenshaw et al., 2009), and an assumed metallicity 5 times solar.
@@
In this model, clouds are launched from the disk surface but remain relatively close to it, with the ratio of vertical height () to radial distance () staying below 5. This limited height is a direct consequence of the low Eddington ratio, giving the BLR a geometry that closely resembles a puffed-up disk surface. The ratio in the model peaks near the inner edge of the BLR, which is defined by the dust sublimation temperature reached at approximately 2260 for the adopted , Eddington rate and metallicity. At this radius, the Keplerian velocity is km s-1 for this , implying that the full width at half maximum (FWHM) of the emission lines could reach up to twice this value. This result based on FRADO model is consistent with the H FWHM of km s-1 reported by Pei et al. (2017). However, contributions from larger radii act to moderate this broadening.
The position of the clouds and their velocity field allow us to construct a 2-D velocity-delay map (see Figure 2, upper panel) which we consider as representative for H line, assuming a viewing angle of degrees, which is likely representative of NGC 5548 (Pancoast et al., 2014; Cappi et al., 2016; Wildy et al., 2021; Horne et al., 2021). Our velocity–delay map closely resembles the predictions of a simple flat Keplerian disk model, as the vertical velocities are minimal, no outflows are included, and no selective shielding is assumed. It shows some similarity in the overall shape to the map derived observationally by Horne et al. (2021) (see their Figures 5 and 7).
However, the main inconsistency lies in the narrower velocity extent observed in our velocity-delay map, which shows emission up to , while Horne et al. (2021) reported observed H emission extending to , although the majority of the response is concentrated within . Note that, FRADO inherently assumes that the BLR originates beyond the dust sublimation radius, which is appropriate for low-ionization lines like H and thus naturally predicts narrower velocity distributions, consistent with emission from outer, slower-moving regions of the BLR. While, the broader velocity distributions observed by Horne et al. (2021) are shaped by contributions from the inner, dust-free high-ionization zones (dominated by CIV, He II lines) of the BLR, where gas orbits closer to the black hole at higher Keplerian speeds.
Nevertheless, the exact comparison is difficult since the observational H map in Horne et al. (2021) seems to be strongly contaminated by He II emission. Notably, while the red wing of the observed H shows a Keplerian structure, the blue wing appears significantly weaker, suggesting substantial He II contamination. In our model, in its present form, we cannot explain such asymmetry. However, such asymmetries in the H line are well-documented and known to evolve over time in a quasi-periodic fashion (Shapovalova et al., 2004). Bon et al. (2016) explained such phenomenon as caused by the presence of the secondary black hole, perturbing BLR, by precession of the outer part of the disk, or spiral waves in the disk. All such mechanisms will give asymmetry between blue and red-shifted wing of the flat BLR coexisting with the disk. In the future we could include the warped disk into FRADO model, but this is beyond the scope of the current paper.
@@
Next, we derive the BLR response function by projecting the map onto time axis. The resulting shape, shown in Figure 2, lower panel, exhibits two distinct peaks: one at shorter time delays, around 2.6 days, corresponding to the clouds on the same side as the observer, and the other peak arising from the clouds on the opposite side of the black hole. The shape of our modeled response function does not perfectly replicate the observed one reported by Horne et al. (2021). In their analysis using the MEMECHO software for H (see Figure 12 for comparison), the projected response exhibits two peaks: a prominent one at approximately 2 days and a second, much smaller peak at around 25 days. While our first peak aligns well with theirs at 2 days, our second peak appears closer in time, at 11 days, and is relatively stronger than the second peak reported by Horne et al. (2021). Despite these differences, it is important to emphasize that our model relies solely on global parameters, without introducing arbitrary values such as inner or outer BLR radii. This absence of free parameters enhances the significance of the observed similarity between the modeled and empirical response functions. Additionally, we assess the consistency between the BLR response function recovered from our FRADO model and the H emission-line response function derived from spectroscopic RM data by Horne et al. (2021) in Section References.
@@
The measured time delay of the H line relative to the 5100 Å continuum during the 2014 campaign (January to June) was reported by Pei et al. (2017) as days, which is shorter than predicted by the radius–luminosity relation. This H delay ranks among the shortest ever recorded for NGC 5548. In contrast, the mean response function from the 19982001 campaign peaked around 20 days (Cackett & Horne, 2006). Horne et al. (2021) later explored the nature of this shorter delay in detail, finding that the response function exhibited a secondary peak at approximately days but began with a high value near to zero time delay. Their observed map also supports the primary response originating from the near side of the BLR, facing the observer.
@@
Additionally, the model predicts an upper limit for the covering factor of the BLR. If the cloud distribution is not transparent, the region intercepts all radiation emitted within inclinations greater than the aspect ratio of the cloud distribution. As shown in Figure 1, the model suggests that less than of the radiation can be intercepted by the BLR, as the cloud height is generally small compared to the radius. This is notably lower than the to covering factor typically expected from BLRs (e.g. Baldwin et al., 1995; Korista & Goad, 2019; Panda, 2021; Panda et al., 2022). In contrast, the yearly variations observed in the line profile of NGC 5548 suggest that the disk can be distorted and/or precessing (e.g., Shapovalova et al., 2009, see their Figure 2). Such a departure from the disk model strictly perpendicular to the symmetry axis may affect the irradiation of the launched clouds and this way to modify the BLR emissivity.
The distribution of BLR clouds enables us to roughly estimate the shape of the H line predicted by the model. In this estimation, we neglect the vertical cloud velocities and only consider the projected rotational velocities towards the observer. The resulting line shape, shown in Figure 3, reveals a two-peak structure that is even more pronounced than that observed in the data. The FWHM predicted by our model approaches 8000 km s-1. In comparison, the measured FWHM of the H line from the rms spectrum was km s-1 (Pei et al., 2017). Notably, the FWHM values have varied over the years, indicating the dynamic nature of the BLR. For instance, Peterson et al. (2004) reported an H FWHM of km s-1, which is in good agreement with our model estimate. These variations suggest that our model captures a plausible state of the BLR, consistent with the historical range of observed line widths.
@@ 4.2 The BLR emissivity
To model the emission characteristics of photoionized clouds, we perform CLOUDY simulations that account for both line and continuum contributions. During our computations, we consider both line and continuum emissions, incorporating radiation from both inward and outward directions to ensure equal visibility of the illuminated and un-illuminated sides of the clouds. Among the most prominent features that emerge are the Balmer and Paschen edges, as illustrated in Figure 4. The depths of these edges are governed by the specific model parameters adopted (see the most recent study by Pandey et al., 2023).
The reprocessing calculations are performed for a single representative cloud, using the full shape of the SED as the incident radiation, as described in Section 3.4. This approach is admittedly a significant simplification, since in reality each cloud is exposed to a different incident spectrum depending on its location and orientation. This spectral variation is properly accounted for in the computation of the radiation force. However, generating comprehensive tables of local emissivity for the full range of incident spectra lies beyond the current scope of this work. Such an extension is essential for future studies to improve the physical realism of the model.